OpenStudy (rayman5766):

The formula d = rt gives the distance d traveled in time t at rate r. A bicyclist rides at a constant rate of 22 kmph. How many kilometers will he travel in 2.5 h?

OpenStudy (rayman5766):

ok wow

OpenStudy (kamille):

d - your distance, r - how fast he drives; t- time he rides so you know r (22) and t (2.5) so d=r*t=22*2.5=?

OpenStudy (rayman5766):

55

OpenStudy (kamille):

correct,little wizard:}